OpenStudy (anonymous):

(yes this a math question but it is closely related to phyiscs) A box is being pulled up a 15 degree incline plane. The force needed is 25 N. Find the horizontal and vertical componets of the force vector and interpret each of them. (the part I need help with is interpreting them)

OpenStudy (bahrom7893):



OpenStudy (bahrom7893):

find x and y using sines and cosines lol

OpenStudy (bahrom7893):

Sin(15) = y/25 Cos(15) = x/25 y is vertical comp, x is horiz comp.

OpenStudy (anonymous):

yes but what do they mean in the problem the numberic values are easy to solve for but interpreting them is hard

OpenStudy (anonymous):

stating one is vertical and one is horizontal is not enough

OpenStudy (bahrom7893):